# Multiclass Local Calibration With the Jensen-Shannon Distance

### v1

@@ Abstract
Developing trustworthy Machine Learning (ML) models requires their predicted probabilities to be well-calibrated, meaning they should reflect true-class frequencies. Among calibration notions in multiclass classification, strong calibration is the most stringent, as it requires all predicted probabilities to be simultaneously calibrated across all classes. However, existing approaches to multiclass calibration lack a notion of distance among inputs, which makes them vulnerable to proximity bias: predictions in sparse regions of the feature space are systematically miscalibrated. This is especially relevant in high-stakes settings, such as healthcare, where the sparse instances are exactly those most at risk of biased treatment. In this work, we address this main shortcoming by introducing a local perspective on multiclass calibration. First, we formally define multiclass local calibration and establish its relationship with strong calibration. Second, we theoretically analyze the pitfalls of existing evaluation metrics when applied to multiclass local calibration. Third, we propose a practical method for enhancing local calibration in Neural Networks, which enforces alignment between predicted probabilities and local estimates of class frequencies using the Jensen-Shannon distance. Finally, we empirically validate our approach against existing multiclass calibration techniques.
@@ 1 Introduction
In many high-stakes applications, Machine Learning (ML) models are expected not only to be accurate but also well-calibrated [25, 21] - i.e., their predicted probabilities should reflect the true empirical frequencies of the corresponding classes. For instance, in the context of clinical decision-making, Van Calster et al. [24] compares two cardiovascular risk prediction models over two million patients from the UK. They show that the better-calibrated model, despite having a slightly lower AUC, provided more useful predictions, avoiding overestimation of risk.
@@
However, existing works have mostly focused on confidence calibration [14, 15, 28, 1], which only looks at the top-predicted class to check a model’s calibration. While in binary classification tasks, a well-calibrated model on the top-predicted class ensures good calibration on the complementary class, this is not true for multiclass classification tasks.
(a) LCN flow
(b) LCN effect on Cifar10 using a resnet-50
@@
Figure 1: Our LoCal Nets (LCN) provide local calibration through feature reshaping (1(a)) Unlike post-hoc calibrators that rescale fixed logits, LCNs jointly (i) learn reduced feature representations $\phi^{\prime}(\mathbf{x})$ and (ii) output new calibrated logits, aligning predictions with local class frequencies via Jensen–Shannon distance. (1(b)) On Cifar10 with resnet-50, LCNs yield tighter, better-separated class clusters and improved calibration ($\approx 64\%$ reduction in MLCE, $36\%$ reduction in LCE).
@@
Example. Consider an ML model trained to predict the stage of cancer in a cell. Suppose the model is calibrated only with respect to its most confident prediction, i.e., if it assigns a probability of $p\%$ to early-stage cancer, then about $p\%$ of the instances receiving such a score are indeed early-stage cancer. Although this might appear sufficient for decision-making [5], inaccurate probability estimates for less frequent classes can be critical: failing to approximate the likelihood of rare transitional states—such as cells halfway between benign and malignant—might hide patterns about tumor progression.
@@
Hence, stronger notions for multiclass calibration have been proposed. One example is strong calibration [23], which enforces alignment across the full probability vector. However, existing multiclass calibration notions do not consider any form of distance among instances, making them prone to proximity bias [28], whereby predictions for instances in sparsely populated regions of the decision space are more likely to be poorly calibrated. In high-stakes settings, these are precisely the cases where one requires trustworthy predictions to avoid biased treatment.
@@
Our Contributions. In this work, we tackle this shortcoming by framing multiclass calibration from a “local” perspective. More precisely, $(i)$ in section 4 we introduce the notion of Multi-Class Local-Calibration, which leverages true empirical frequencies in the neighborhoods of inputs to assess model reliability across the decision-space, and we connect it to strong calibration; $(ii)$ we provide theoretical insights on existing pitfalls of current evaluation metrics when applied to multiclass local calibration (Sections 56); $(iii)$ in section 7 we propose a new neural network calibration method called LoCal Nets (LCNs), which exploits the Jensen-Shannon distance to align predicted probabilities and local estimates of class frequencies, while enforcing denser feature representations (Figure 1); $(iv)$ we evaluate our approach against existing competitors (section 8), showing how our method improves over local calibration metrics and keeps competitive performance at the global level.
@@ 2 Related Work
Global Calibration Notions.
Due to the stringent requirements of strong calibration and difficulty in both evaluation and enforcement, the literature has proposed several relaxed alternatives. Examples of such relaxations are confidence calibration [5] and top-label calibration [6]. While the former considers only the model’s maximum predicted confidence, the latter refines the definition of confidence calibration by conditioning the empirical frequencies on both the model confidence and the model prediction.
Class-wise calibration
@@
[13] evaluates calibration for each class marginally, by comparing predicted probabilities with empirical frequencies for that target alone. Top-$r$ calibration [7] considers if the true class falls within the top-$r$ predicted labels and if the cumulative confidence over the top-$r$ classes aligns with observed frequencies. Decision calibration [32] defines calibration with respect to a decision-making policy, requiring predicted and empirical distributions to match according to a decision maker. Since these notions are global, they emphasize predicted probabilities while overlooking factors such as instance density or spatial position, failing to capture poor calibration for underrepresented groups.
Post-hoc Calibration Methods.
@@
When a trained model suffers from poorly calibrated outputs, many available model-agnostic post-hoc techniques can be employed to adjust its predicted probabilities. Popular methods include binning [30], isotonic regression [31], logistic scaling [20], temperature scaling [5] and parametric forms of scaling [12, 13]. We refer to de Menezes e Silva Filho et al. [1] for a complete overview of different calibration techniques. However, all these methods treat logits as fixed inputs to be rescaled. In contrast, our approach learns new feature representations that reshape the geometry of the representation space itself. Thus, these new neighborhoods better reflect true class frequencies and improve local calibration.
Local Calibration Notions.
@@
Most methods group instances with similar predicted confidences, but ignore that similar scores can be assigned to points in very different regions of the decision space. A first attempt to investigate the notion of local calibration is by Luo et al. [15], who propose to recalibrate a classifier’s confidence scores via kernel regressions. Although promising, their approach only tackles confidence calibration and considers fixed feature representations. As a consequence, the method significantly slows down inference times, requiring a kernel regression for each recalibrated probability score. Also [16] investigates the utility of kernel regressions for calibration, but their proposal directly applies to a model training process. As such, this approach can suffer from computational and data limitations. Finally,  Perez-Lebel et al. [19] shows that whenever the classifier’s decision boundary is complex, it can lead to poor calibration of the scores predicted for less likely instances. They introduce (i) the cancellation effect, where miscalibration errors within a confidence group offset one another, and (ii) proximity bias, reflecting disparities in calibration quality for instances in sparsely populated regions of the decision space. Our approach differs as they focus on (i) data density rather than on the position of points in the decision space; (ii) their proposal addresses only the confidence score of the predicted class, not multiclass calibration. We refer readers to the Appendix C for a detailed analysis of the role of proximity bias in our proposal and an illustrative example of the dangers of density-based recalibration.
@@ 3 Background
Let us consider a multi-class classification setting, where $\mathcal{X}\subseteq\mathbb{R}^{m}$ is the feature space and $\mathcal{Y}=\{0,\ldots,C-1\}$ is a finite target space with $C$ distinct labels. Let us assume we have access to a given dataset $D=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$ of input-output pairs drawn from an unknown joint distribution $\mathcal{P}$ over $\mathcal{X}\times\mathcal{Y}$. Each input $\mathbf{x}_{i}\in\mathcal{X}$ is a feature vector of $m$ dimensions, and each label $y_{i}\in\mathcal{Y}$ has a corresponding one-hot encoded vector $\mathbf{y}_{i}$ indicating the correct class among the $C$ possible classes. We consider a probabilistic classifier $f\colon\mathcal{X}\rightarrow\Delta^{C}$, where $\Delta^{C}$ is the (C1)(C-1)-dimensional probability simplex. In words, a probabilistic classifier maps an input $\mathbf{x}$ to a probability distribution over classes, i.e., $f(\mathbf{x})=\mathbf{\hat{p}}\in\Delta^{C}$, where each entry $\mathbf{\hat{p}}_{k}=f_{k}(\mathbf{x})$ of the predicted probability vector $\mathbf{\hat{p}}$ denotes the predicted probability of class $k$.
In the multiclass context, the weakest notion of calibration is confidence calibration [5], which requires that only the top-predicted probability $\max_{y\in\mathcal{Y}}\hat{\mathbf{p}}_{y}$ matches the frequency of correct predictions of the classifier. A more stringent notion is strong calibration [23], which requires the target class conditional distribution on any prediction of the classifier to match that prediction, i.e.,:
@@
Metrics.
Because of the large number of calibration notions, various metrics have been proposed in the literature to evaluate the calibration of classifiers.
One of the most popular metrics is Expected Calibration Error ($ECE$) [18], which measures the calibration of a binary classifier. It bins predicted confidence scores into $B$ intervals and compares the average predicted confidence with the empirical accuracy in each bin. A multiclass extension of $ECE$ is Class-wise ECE, which calculates ECE separately for each class $c$ and then averages the results. Multidimensional Expected Calibration Error ($MECE$) extends $ECE$ and its class-wise version to multi-class, by using a multidimensional grid binning. Unfortunately, the combinatorial nature of such a binning hinders its practical application. Expected Cumulative Calibration Error (ECCE) [9] evaluates the cumulative discrepancy between confidence and accuracy across bins rather than averaging them.
Local Calibration Error ($LCE$) quantifies directly the state of local calibration of a probabilistic classifier. We provide here the multiclass extension from the original definition by Luo et al. [15]:
@@
where $\mid\mid\cdot\mid\mid_{1}$ is an appropriate $\ell^{1}$ norm, $m_{b}$ the number of used bins, $B_{b}$ is the set of instances in the $b$-th bin and $k_{\gamma}(\mathbf{x}_{i},\mathbf{x}_{j})$ is a kernel function that weights the influence of neighbouring points of the anchor $\mathbf{x}_{i}$ to its individual $LCE$ score. Its appearance in the denominator as a normalization term ensures weights sum to 1 for every anchor. In practice, this metric captures the differences in the predicted probabilities and the corresponding ground truths for neighbours of an anchor point $\mathbf{x}_{i}$. Finally, the Maximum Local Calibration Error is $MLCE=\max_{i\in D}\left\|\frac{\sum_{j\in b}\bigl(\hat{\mathbf{p}}_{j}-\mathbf{y}_{j}\bigr)\,k_{\gamma}(\mathbf{x}_{i},\mathbf{x}_{j})}{\sum_{j\in b}k_{\gamma}(\mathbf{x}_{i},\mathbf{x}_{j})}\right\|_{1}.$ Notice that, in its original formulation by Luo et al. [15], $LCE$ is limited to confidence calibration, making it insufficient to assess multiclass local calibration.
@@ 4 Defining Multiclass Local Calibration
Let $f$ be a probabilistic classifier, and let $D$ be an evaluation set drawn from $\mathcal{P}$. For each instance $i\in D$, let $k_{\gamma}(\mathbf{x}_{i},\mathbf{x}_{j})$, with $j\neq i\in D$ be a kernel function with bandwidth $\gamma$ and consider an associated kernel estimator $\hat{\theta}(\mathbf{y}_{i}\mid\mathbf{x}_{i})=\frac{\sum_{j\in D}W_{j}(\mathbf{x}_{i})\mathbf{y}_{j}}{\sum_{j\in D}W_{j}(\mathbf{x}_{i})}$ with $W_{j}(\mathbf{x}_{i})\propto k_{\gamma}(\mathbf{x}_{i},\mathbf{x}_{j})$, a kernel induced weight, normalized to sum to $1$. Finally, let $\hat{\theta}(\mathbf{y}_{i}\mid\mathbf{x}_{i})$ be consistent in the mean squared error (MSE) sense.
Then, $f$ is locally calibrated on $D$ if, for all $\mathbf{x}_{i}\in D$, the predicted probability vector $\hat{\mathbf{p}}_{i}=f(\mathbf{x}_{i})$ is close to the kernel estimate up to a tolerance $\varepsilon\geq 0$, i.e.,
$\left\|\mathbf{\hat{p}}_{i}-\hat{\theta}(\mathbf{y}_{i}\mid\mathbf{x}_{i})\right\|_{1}\leq\varepsilon,\quad\forall i\in\{1,\dots,n\}.$ (2)
@@
Notably, the multiclass notion of local calibration is closely related to strong calibration, and satisfying the former is sufficient for the latter. In the theorem below, we establish a bound in terms of the continuous $MECE$. We focus on the continuous form since, unlike binned variants, it integrates over the probability space and thus provides the most faithful and comprehensive measure of calibration.
@@
Let $D$ be an evaluation dataset drawn i.i.d. from a distribution $\mathcal{P}$. Define the continuous Multidimensional Expected Calibration Error (MECE) as:
$M(f)=\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x})}\mathbb{E}_{y\sim p(y)}\left[\left|\mathbb{E}[\mathbf{y}\mid\mathbf{x}]_{y}-f_{y}(\mathbf{x})\right|\right],$
If a model $f$ satisfies local calibration, then there exists $k\in[1/C,1]$ such that continuous $MECE$ is asymptotically upper bounded as:
@@
Theorem 3 shows that the continuous $MECE$ is upper bounded for multiclass locally calibrated models and is minimal if the condition holds for small $\varepsilon$. A detailed proof is provided in the Appendix B.

### v2

@@ Abstract
Developing trustworthy Machine Learning (ML) models requires their predicted probabilities to be well-calibrated, meaning they should reflect true-class frequencies. Among calibration notions in multiclass classification, strong calibration is the most stringent, as it requires all predicted probabilities to be simultaneously calibrated across all classes. However, existing approaches to multiclass calibration lack a notion of distance among inputs, which makes them vulnerable to proximity bias: predictions in sparse regions of the feature space are systematically miscalibrated. In this work, we address this main shortcoming by introducing a local perspective on multiclass calibration. First, we formally define multiclass local calibration and establish its relationship with strong calibration. Second, we theoretically analyze the pitfalls of existing evaluation metrics when applied to multiclass local calibration. Third, we propose a practical method to enhance local calibration in Neural Networks, which enforces alignment between predicted probabilities and local estimates of class frequencies using the Jensen-Shannon distance. Finally, we empirically validate our approach against existing multiclass calibration techniques.
Multiclass Local Calibration with the Jensen-Shannon Distance
Cesare Barbera⋄,∗,†          Lorenzo Perini          Giovanni De Toni
Andrea Passerini          Andrea Pugnana⋄,∗
University of Trento          University of Pisa          Meta          Fondazione Bruno Kessler
@@ 1 INTRODUCTION
00footnotetext: Corresponding authors. {name}.{surname}@unitn.it.
@@
In many high-stakes applications, Machine Learning (ML) models are expected not only to be accurate but also well-calibrated (Wang, 2023; Sambyal et al., 2023; Cianci et al., 2023; Ruggieri and Pugnana, 2025) - i.e., their predicted probabilities should reflect the true empirical frequencies of the corresponding classes. For instance, in the healthcare context, Van Calster et al. (2019) compares two cardiovascular risk prediction models over two million patients from the UK. They show that the better-calibrated model, despite having a lower AUC, avoided overestimation of risk.
@@
However, existing works have mostly focused on confidence calibration (Le-Coz et al., 2024; Luo et al., 2022; Xiong et al., 2023; de Menezes e Silva Filho et al., 2023), which only looks at the top-predicted class to check a model’s calibration. While in binary classification tasks, a well-calibrated model on the top-predicted class ensures good calibration, this is not true for multiclass classification tasks.
(a) LN flow
(b) LN effect on Cifar10 using a resnet-50
@@
Figure 1: Our LoCal Nets (LN) provide local calibration through feature reshaping (1(a)) Unlike post-hoc calibrators that rescale frozen ( Refer to caption ) logits, LNs jointly (i) learn reduced feature representations $\phi^{\prime}(\mathbf{x})$ and (ii) output new calibrated logits, aligning predictions with local class frequencies via Jensen–Shannon distance. (1(b)) On Cifar10 with resnet-50, LNs yield tighter, better-separated class clusters and improved calibration ($\approx 64\%$ reduction in MLCE, $36\%$ reduction in LCE).
@@
Example. Consider an ML model trained to predict the stage of cancer in a cell. Suppose the model is calibrated only with respect to its most confident prediction, i.e., if it assigns a probability of $p\%$ to early-stage cancer, then about $p\%$ of the instances receiving such a score are indeed early-stage cancer. Although this might appear sufficient for decision-making (Guo et al., 2017), inaccurate probability estimates for less frequent classes can be critical: failing to approximate the likelihood of rare transitional states—such as cells halfway between benign and malignant—might hide patterns about tumor progression.
@@
Hence, stronger notions for multiclass calibration have been proposed. One example is strong calibration (Vaicenavicius et al., 2019), which requires alignment across the full probability vector. However, existing multiclass calibration notions do not consider distances among instances, making them prone to proximity bias (Xiong et al., 2023), whereby predictions for instances in sparsely populated regions of the decision space are more likely to be poorly calibrated. This reveals a structural limitation of existing multiclass calibration notions and motivates the need for definitions that explicitly incorporate geometric structure.
@@
Our Contributions. In this work, we tackle this shortcoming by framing multiclass calibration from a “local” perspective. More precisely, $(i)$ in Section 4 we introduce the notion of Multi-Class Local-Calibration, which leverages true empirical frequencies in the neighborhoods of inputs to assess model reliability across the decision-space, and we connect it to strong calibration; $(ii)$ we provide theoretical insights on existing pitfalls of current evaluation metrics when applied to multiclass local calibration (Sections 56); $(iii)$ in Section 7 we propose a new neural network calibration method called LoCal Nets (LNs), which exploits the Jensen-Shannon distance to align predicted probabilities and local estimates of class frequencies, while enforcing denser feature representations (Figure 1); $(iv)$ we evaluate our approach against existing competitors (Section 8), showing how our method improves over local calibration metrics and keeps competitive performance at the global level.
@@ 2 RELATED WORK
Global Calibration Notions. Due to the stringent requirements of strong calibration and difficulty in both evaluation and enforcement, the literature has proposed several relaxed alternatives. Examples of such relaxations are confidence calibration (Guo et al., 2017) and top-label calibration (Gupta and Ramdas, 2022). While the former considers only the model’s maximum predicted confidence, the latter refines the definition of confidence calibration by conditioning the empirical frequencies on both the model confidence and the model prediction. Class-wise calibration (Kull et al., 2019) evaluates calibration for each class marginally, by comparing predicted probabilities with empirical frequencies for that target alone. Top-$r$ calibration (Gupta et al., 2021) considers if the true class falls within the top-$r$ predicted labels and if the cumulative confidence over the top-$r$ classes aligns with observed frequencies. Decision calibration (Zhao et al., 2021) defines calibration with respect to a decision-making policy, requiring predicted and empirical distributions to match according to a decision maker. Since these notions are global, they emphasize predicted probabilities while overlooking factors such as instance density or spatial position, failing to capture poor calibration for underrepresented groups.
@@
Post-hoc Calibration Methods. When a trained model suffers from poorly calibrated outputs, many available model-agnostic post-hoc techniques can be employed to adjust its predicted probabilities. Popular methods include binning (Zadrozny and Elkan, 2001), isotonic regression (Zadrozny and Elkan, 2002), logistic scaling (Platt and others, 1999), temperature scaling (Guo et al., 2017) and parametric forms of scaling (Kull et al., 2017, 2019; Berta et al., 2026). We refer to de Menezes e Silva Filho et al. (2023) for a complete overview of different calibration techniques. However, all these methods treat logits as fixed inputs to be rescaled. In contrast, our approach learns new feature representations that reshape the geometry of the representation space itself. Thus, these new neighborhoods better reflect true class frequencies and improve local calibration.
@@
Local Calibration Notions. Most methods group instances with similar predicted confidences, but ignore that similar scores can be assigned to points in very different regions of the decision space. Hence, researchers focused on approaches that incorporate a notion of distance between inputs. Luo et al. (2022), Lin et al. (2023), and Valk and Kull (2023) propose recalibration approaches based on kernel regression. Luo et al. (2022) focus on confidence calibration, leveraging kernel smoothing to locally adjust predicted confidences. In contrast, Lin et al. (2023) and Valk and Kull (2023) extend kernel-based recalibration to the multiclass setting. A key distinction concerns how feature representations are treated. Both Luo et al. (2022) and Valk and Kull (2023) operate on fixed representations. In contrast, Lin et al. (2023) learn new feature representations tailored to kernel-based estimates. Our approach is closest in spirit to the latter, yet it differs fundamentally in its objective. Rather than using kernel regression as the final predictor, we use kernel-based local frequency estimates only as a supervision signal to induce locality in the model’s probabilistic output. Moreover, these approaches do not explicitly formalize or evaluate multiclass calibration as a local property. In contrast, our work introduces a formal notion of multiclass local calibration and directly studies its theoretical and empirical implications. Also Marx et al. (2023) investigates the utility of kernel regressions for calibration, but their proposal directly applies to a model training process. As such, this approach can suffer from computational and data limitations. Finally,  Perez-Lebel et al. (2023) shows that whenever the classifier’s decision boundary is complex, it can lead to poor calibration of the scores predicted for less likely instances. They introduce (i) the cancellation effect, where miscalibration errors within a confidence group offset one another, and (ii) proximity bias, reflecting disparities in calibration quality for instances in sparsely populated regions of the decision space. Our approach differs as they focus on (i) data density rather than on the position of points in the decision space; (ii) their proposal only addresses the confidence score of the predicted class. We provide in Section C.2 a detailed analysis of the role of proximity bias in our proposal and an illustrative example of the pitfalls of density-based recalibration.
Multicalibration. Another notion, known as Multicalibration, requires the model to be calibrated not just on the entire population, but on virtually any sub-population identifiable by a specific hypothesis class (Hébert-Johnson et al., 2018; Jung et al., 2021; Baldeschi et al., 2025; Jin et al., 2025; Tax et al., 2026). Thus, the goal of multicalibration approaches is to ensure that models do not make biased predictions on any of these sub-populations (Globus-Harris et al., 2023; Noarov and Roth, 2023).
While related, multicalibration differs from our setting in two key aspects. First, it is mostly studied for binary/regression tasks, whereas we address multiclass calibration and the complexities of the probability simplex. Second, it identifies semantically meaningful subgroups (e.g., User Age > 18) using expressive features in tabular data. In contrast, our approach operates in the high-dimensional representation space of neural networks, where dimensions are abstract and “auditing” for subgroups via standard multicalibration techniques is computationally intractable.
@@ 3 BACKGROUND
In this section, we formally introduce the main definitions and metrics used to evaluate calibration.
Definitions. Let us consider a multi-class classification setting, where $\mathcal{X}\subseteq\mathbb{R}^{m}$ is the feature space and $\mathcal{Y}=\{0,\ldots,C-1\}$ is a finite target space with $C$ distinct labels. Let us assume we have access to a given dataset $D=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$ of input-output pairs drawn from an unknown joint distribution $\mathcal{P}$ over $\mathcal{X}\times\mathcal{Y}$. Each input $\mathbf{x}_{i}\in\mathcal{X}$ is a feature vector of $m$ dimensions, and each label $y_{i}\in\mathcal{Y}$ has a corresponding one-hot encoded vector $\mathbf{y}_{i}$ indicating the correct class among the $C$ possible classes. We consider a probabilistic classifier $f\colon\mathcal{X}\rightarrow\Delta^{C}$, where $\Delta^{C}$ is the (C1)(C-1)-dimensional probability simplex. In words, a probabilistic classifier maps an input $\mathbf{x}$ to a probability distribution over classes, i.e., $f(\mathbf{x})=\mathbf{\hat{p}}\in\Delta^{C}$, where each entry $\mathbf{\hat{p}}_{k}=f_{k}(\mathbf{x})$ of the predicted probability vector $\mathbf{\hat{p}}$ denotes the predicted probability of class $k$.
In the multiclass context, the weakest notion of calibration is confidence calibration (Guo et al., 2017), which requires that only the top-predicted probability $\max_{y\in\mathcal{Y}}\hat{\mathbf{p}}_{y}$ matches the frequency of correct predictions of the classifier. A more stringent notion is strong calibration (Vaicenavicius et al., 2019), which requires the target class conditional distribution on any prediction of the classifier to match that prediction, i.e.,:
@@
Metrics. Because of the large number of calibration notions, various metrics have been proposed in the literature to evaluate the calibration of classifiers.
One of the most popular metrics is Expected Calibration Error ($ECE$) (Naeini et al., 2015), which measures the calibration of a binary classifier. It bins predicted confidence scores into $B$ intervals and compares the average predicted confidence with the empirical accuracy in each bin. A multiclass extension of $ECE$ is Class-wise ECE, which calculates ECE separately for each class $c$ and then averages the results. Multidimensional Expected Calibration Error ($MECE$) extends $ECE$ and its class-wise version to multi-class, by using a multidimensional grid binning. Unfortunately, the combinatorial nature of such a binning hinders its practical application. Expected Cumulative Calibration Error (ECCE) (Ibarra et al., 2022) evaluates the cumulative discrepancy between confidence and accuracy across bins rather than averaging them.
Local Calibration Error ($LCE$) quantifies directly the state of local calibration of a probabilistic classifier. We provide here the multiclass extension from the original definition by Luo et al. (2022):
@@
where $\mid\mid\cdot\mid\mid_{1}$ is an appropriate $\ell^{1}$ norm, $m_{b}$ the number of used bins, $B_{b}$ is the set of instances in the $b$-th bin and $k_{\gamma}(\mathbf{x}_{i},\mathbf{x}_{j})$ is a kernel function that weights the influence of neighbouring points of the anchor $\mathbf{x}_{i}$ to its individual $LCE$ score. Its appearance in the denominator as a normalization term ensures weights sum to 1 for every anchor. In practice, this metric captures the differences in the predicted probabilities and the corresponding ground truths for neighbors of an anchor point $\mathbf{x}_{i}$. The shown formalization assumes balanced classes for simplicity and is presented in the most general form leveraging multidimensional binning. In practice, a trivial class-wise adaptation can be implemented. Finally, the Maximum Local Calibration Error is $MLCE=\max_{i\in D}\left\|\frac{\sum_{j\in b}\bigl(\hat{\mathbf{p}}_{j}-\mathbf{y}_{j}\bigr)\,k_{\gamma}(\mathbf{x}_{i},\mathbf{x}_{j})}{\sum_{j\in b}k_{\gamma}(\mathbf{x}_{i},\mathbf{x}_{j})}\right\|_{1}.$ Notice that, in its original formulation by Luo et al. (2022), $LCE$ is limited to confidence calibration, making it insufficient to assess multiclass local calibration.
@@ 4 DEFINING MULTICLASS LOCAL CALIBRATION
For each instance $i\in D$, let $k_{\gamma}(\mathbf{x}_{i},\mathbf{x}_{j})$, with $j\neq i\in D$ be a kernel function with bandwidth $\gamma$. We consider bandwidth sequences $\{\gamma_{n}\}$ satisfying standard consistency conditions. Define the associated kernel estimator:
$\hat{\theta}(\mathbf{y}_{i}\mid\mathbf{x}_{i})=\frac{\sum_{j\in D}W_{j}(\mathbf{x}_{i})\mathbf{y}_{j}}{\sum_{j\in D}W_{j}(\mathbf{x}_{i})},\;\quad\;W_{j}(\mathbf{x}_{i})\propto k_{\gamma}(\mathbf{x}_{i},\mathbf{x}_{j}),$ (2)
where the weights are normalized to sum to $1$. Then, $f$ is locally calibrated on $D$ if, for all $\mathbf{x}_{i}\in D$, the predicted probability vector $\hat{\mathbf{p}}_{i}=f(\mathbf{x}_{i})$ is close to the kernel estimate up to a tolerance $\varepsilon\geq 0$, i.e.,
$\left\|\mathbf{\hat{p}}_{i}-\hat{\theta}(\mathbf{y}_{i}\mid\mathbf{x}_{i})\right\|_{1}\leq\varepsilon,\quad\forall i\in\{1,\dots,n\}$ (3)
@@
Please refer to Section A.2 for a comprehensive description of the consistency conditions. We stress that Definition 1 is inherently finite-sample and bandwidth-dependent as it requires alignment with the statistically optimal local frequency estimate at the chosen resolution $\gamma$, not with the true conditional distribution. The consistency assumptions on the kernel estimator relate this empirical notion to population-level calibration measures. Notably, the notion of multiclass local calibration is structurally related to strong calibration. On the one hand, strong calibration enforces the alignment of predicted probabilities with empirical frequencies conditional on the predicted probability vector. On the other hand, local calibration constrains this alignment to be pointwise in the feature space. As we show in Eq. 4, satisfying local calibration is asymptotically sufficient to guarantee strong calibration. More precisely, we establish an upper bound on the continuous Multidimensional Expected Calibration Error (MECE). We focus on this formulation since, unlike its binned counterparts, it integrates over the probability simplex and provides a distribution-level characterization of calibration.
@@
Define the continuous Multidimensional Expected Calibration Error (MECE) as:
$M(f)=\sum_{y\in\mathcal{Y}}p(y)\,\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x})}\left[\left|\mathbb{E}[\mathbf{y}\mid\mathbf{x}]_{y}-f_{y}(\mathbf{x})\right|\right],$
If a model $f$ satisfies local calibration, under standard regularity conditions there exists $k\in(0,1]$ such that continuous $MECE$ is asymptotically upper bounded:
@@
Thus, in the limit, local calibration can be interpreted as a geometric refinement of strong calibration, ensuring that global reliability emerges from consistent local behavior. A detailed proof is in Section B.1.